OpenStudy (anonymous):

A certain mountain has an elevation of 19,517ft. In 1916, the glacier on this peak covered 4 acres. By 2004 this glacier had melted to only 1 acre. a) Assume that this glacier melted at a constant rate each year. Find the yearly rate. b) Use the anser from part (a) to write a linear equation that gives the acreage A of this glacier t years past 1916.

OpenStudy (anonymous):

i think whne they ask for the yearky rate it meanst the slope

OpenStudy (anonymous):

m= (y2-y1)/(x2-x1) m= (1-4)/(2004-1916)

OpenStudy (anonymous):

m=-3/88 for the b part we use the equation for a line y-y0=m(x-x0) y-4=(-3/88)(x-0) asumming that 1916 is the year 0, y= adreage A, x= t (time)